# Numaan NaeemKaushal Kumar MauryaKseniia PetukhovaEkaterina Kochmar

### v1

@@ Abstract
Numaan Naeemthanks: Equal contribution.   Kaushal Kumar Maurya   Kseniia Petukhova   Ekaterina Kochmar Mohamed bin Zayed University of Artificial Intelligence (MBZUAI) {Numaan.Naeem,Kaushal.Maurya,Kseniia.Petukhova,Ekaterina.Kochmar}@mbzuai.ac.ae
@@ 1 Introduction
The current version of our tool focuses on the pedagogical quality evaluation of tutor responses in the context of Student Mistake Remediation (SMR) Boaler (2013); Handa et al. (2023) in the mathematical domain, where tutors address errors or misconceptions that hinder student progress Wang et al. (2024a, b); Macina et al. (2023a). As the foundation for evaluation, we use the established taxonomy from Maurya et al. (2025), which is grounded in the learning sciences principles and which allows us to assess the quality of tutor responses along four key SMR dimensions Kochmar et al. (2025): (1) mistake identification, concerned with whether tutor’s response notifies the student of the committed mistake; (2) mistake location, focusing on whether the tutor clearly points to the erroneous part in the student’s solution; (3) providing guidance, evaluating the quality of the pedagogical guidance; and (4) actionability, assessing whether tutor’s response makes it clear what the student should do next. These dimensions are further outlined in Table 2, with an illustrative example provided in Figure 1.
The structure and implementation of the front- and backend of our tool are described in Section 3. Using MRBench dataset Maurya et al. (2025) and taking inspiration from the BEA 2025 shared task on AI tutor response evaluation Kochmar et al. (2025), we introduce a novel, efficient, and light-weight multi-task learning model that addresses the four dimensions of pedagogical quality evaluation ($\S$ 3.1). The outputs of the model, as well as those of an open-source (Prometeus Kim et al. (2024)) and a commercial (GPT-5) LLMs used as judges, are displayed using an interactive browser-based UI with helpful visualizations ($\S$ 3.2). The evaluation results (Section 4) suggest that while our model achieves competitive results when evaluated against gold standard annotation, its outputs are also perceived by users to be at least as accurate as those of the commercial LLM-as-judge models, and the UI is considered informative and easy to use. We present and publicly release:
@@
  • An interactive UI available at https://demo-ai-tutor.vercel.app, which communicates the results and showcases the capabilities of AI tutors in an interpretable way, which we consider to be of interest to education stakeholders and the *ACL community at large. The demo tool can also be run locally with the user’s own data and models.
@@ 2 Related Work
Over the years, the NLP community has seen significant advances in the development of publicly available toolkits for modeling and evaluation, including Hugging Face Transformers Wolf (2019), NLTK Bird (2006), and Scikit-learn Pedregosa et al. (2011). These toolkits have enhanced code reusability, enabling researchers to focus on developing more sophisticated models and metrics. However, in the educational domain, especially in conversational dialogues, there remains a lack of robust tools to push research boundaries. Some progress has been made with frameworks like ConvoKit Chang et al. (2020), which facilitates the analysis of general conversational data. More recently, Edu-ConvoKit Wang and Demszky (2024) was developed for preprocessing, annotation, and analysis of educational dialogues. While these toolkits contribute significantly to data handling and analysis, they fall short of addressing the evaluation of pedagogical quality in AI-driven educational systems. To the best of our knowledge, there is no toolkit that focuses on the on-the-fly assessment of the pedagogical abilities of AI tutors. With AITutor-EvalKit, we aim to fill this critical gap by providing an open-source toolkit that integrates popular taxonomies like that of Maurya et al. (2025) into a single unified framework for the systematic evaluation of AI tutors and can be extended with further approaches such as that of Macina et al. (2025), thereby fostering progress in this underexplored yet important research area.
Figure 2: AITutor-EvalKit pipeline: The backend module includes several model options to assess the pedagogical soundness of tutors’ responses, and the frontend projects evaluation outputs via an interactive user interface.
@@ 3 System and Demonstration Description
AITutor-EvalKit consists of two major modules: backend and frontend, with the pipeline illustrated in Figure 2. The backend module includes several models to evaluate tutor responses, and frontend seamlessly integrates these evaluation outputs with an interactive, customizable, and flexible UI. Different audience groups can benefit from different functionalities of the toolkit: e.g., researchers and developers can use both modules, train their own automated models, use their own datasets, and launch the demo UI locally, while teachers and policymakers can use the frontend module to understand the capabilities of LLMs-as-tutors and make decisions accordingly.
@@ 3.1 Backend Module: Evaluation Models
The backend module has two components: automated evaluation and evaluation by LLM-as-a-judge. In this section, we provide details on the training data (for the automated evaluation model), test data, and automated evaluation model itself, as well as the usage of open-source LLMs and closed-source GPT models as judges.
@@ 3.1.1 Dataset
As discussed in Section 1, we used MRBench dataset by Maurya et al. (2025), which has 491 dialogues (300 in the development set and 191 in the test set), each paired with seven LLM-generated tutor responses and two human tutor responses. Each response is annotated by human annotators for each of the four dimensions using the categories “Yes,” “To some extent,” and “No.” We also created a randomly selected demonstration set consisting of 10 dialogues, which is a subset of the test set. More details on the dimensions, annotation, and data statistics can be found in Appendix A.
3.1.2 Automated Evaluation: LoMTL
This component implements a ternary classifier to evaluate tutor responses across the four considered dimensions. The models developed by participants in the BEA shared task provide a good starting point, as summarized by Kochmar et al. (2025). However, upon closer inspection, we found that several teams relied on closed-source models, fully fine-tuned models, LoRA-tuned models, and some even used multiple models or ensembles for each dimension, which makes these approaches difficult to scale across all dimensions.
@@ 3.1.3 LLM-as-a-Judge Evaluation
This component provides functionality for evaluation using LLMs-as-judges. Following Maurya et al. (2025), we have selected Prometheus2 Kim et al. (2024) as the primary open-source LLM. However, the implementation of this module is flexible enough to support the family of Llama and other open-source causal LLMs for evaluation. Additionally, we have included an option to evaluate using the closed-source OpenAI GPT-5 model, although the implementation supports any model from the GPT family. Users must provide their OpenAI API key to run evaluations.
@@ 3.2 Frontend Module: Demo App
We present an interactive prototype built on top of MRBench, designed to evaluate the pedagogical abilities of AI tutors in educational dialogues. Users can explore the demo in two modes: (1) a Static Mode, where educators can access a deployed version containing 10 conversations from the MRBench test set evaluated by our fine-tuned model and share feedback on the helpfulness of tutor responses; and (2) a Customized Mode, where developers can run the interface locally to analyze their own datasets using either our LoMTL model or their own one, supported by the provided code. The UI consists of three key modules – Automated Evaluation, LLM Evaluation, and Visualizer, that together enable exploration, analysis, and visualization of educational dialogues.
@@ 3.2.1 Automated Evaluation UI
The Automated Evaluation module presents results generated by our LoMTL model, assessing AI tutors across four pedagogical dimensions. As shown in Figure 6, it begins with an overview panel summarizing key statistics such as the number of topics, models, dimensions, and conversations. Users can evaluate a single tutor or compare two tutors side by side.
Single Tutor Evaluation: In this mode, users select a problem topic to view the complete student-tutor dialogue in the “Context” block (Figure 7). A drop-down menu allows selection of a tutor model, and the corresponding response is displayed in the “Tutor Response” block. Users can rate the usefulness of the response (Helpful, To Some Extent, or Not Helpful) and optionally view the ground truth solution for better context. Upon clicking “Get Auto-Evaluation Results,” the system generates performance evaluations across the chosen dimensions, with results downloadable in PNG, JPG, or JSON format. A “Best Performance by Dimension” panel highlights the top tutor(s) for each dimension, helping users quickly identify their pedagogical strengths.
@@ 3.2.2 LLM Evaluation UI
The LLM Evaluation module enables advanced pedagogical assessment of AI tutor responses using LLMs as judges. It extends the functionality of the automated evaluation pipeline by leveraging LLMs to judge and compare tutor responses across four pedagogical dimensions. The UI supports three evaluation modes: single tutor evaluation, tutor model comparison, and judge LLM comparison. As shown in Figure 10, the overview panel summarizes available topics, conversations, tutor models, evaluation dimensions, and judge LLMs, providing a quick snapshot of the evaluation setup. Currently, two LLM judges are supported: GPT-5 and Prometheus-7B-v2.0 Kim et al. (2024).
@@ 3.2.3 Visualizer UI
The Visualizer module provides a high-level overview of the MRBench development set across four evaluation dimensions through intuitive visual analytics. It begins with a “Dataset Overview” panel summarizing key statistics, including the number of conversations, tutor models, and evaluation dimensions. The module includes three main visualization panels: Tutor Performance Summary, Visualization Controls, and Dataset Visualization.
The Tutor Performance Summary panel presents average scores for each tutor model across all four dimensions, where “Yes,” “To some extent,” and “No” correspond to 1.0, 0.5, and 0.0, respectively (Figure 12). The Visualization Controls panel allows users to select specific tutors and dimensions to generate detailed visualizations (Figure 13). The Dataset Visualization panel then displays the results through spider and bar charts, where spider plots summarize tutor strengths and weaknesses across dimensions (Figure 14), while bar charts show detailed score distributions for selected dimensions (Figure 15), along with averages for each response label.
@@ 4 Evaluation
To evaluate the system, we first measure its performance using the metrics adopted from the BEA 2025 shared task Kochmar et al. (2025) – accuracy and macro-F1, and then conduct a human evaluation study, in which participants evaluate both the prediction quality of the LoMTL evaluation model and the usability of our demo tool and UI.
@@ 4.1 Intrinsic Evaluation: Quantitative Analysis
For intrinsic evaluation, we compare our LoMTL model’s predictions on the test set from Kochmar et al. (2025) with the gold human annotations and also run GPT-5 and Prometheus2 on the same data. The average results across evaluation dimensions, presented in Table˜1, show that Prometheus2 performs substantially worse than both our model and GPT-5. Our model achieves the highest averaged accuracy and macro-F1 on the full test set, and it also performs competitively on the demonstration subset. While GPT-5 achieves a slightly higher averaged macro-F1 on the demonstration subset, the overall trend indicates that our model provides more reliable evaluations than Prometheus2 and performs on par with, or better than, GPT-5. The per-dimension results are provided in Table˜6 of Appendix C.1.
@@
Prometheus2 0.47 0.41 0.41 0.33
GPT-5 0.66 0.57 0.66 0.58
Table 1: Accuracy and macro-F1 scores (averaged across Mistake Identification (MI), Mistake Location (ML), Providing Guidance (PG), and Actionability (AC)) for our LoMTL model, Prometheus2, and GPT-5 on the full test set from Kochmar et al. (2025) and on the demonstration set. Best results are shown in bold.
@@ 4.2 Extrinsic Evaluation: Human Study
5 Future Work
Future work will extend the toolkit by (1) integrating new user dialogues and LLMs in the frontend module to generate real-time responses and evaluations; (2) expanding the evaluation dimensions; and (3) broadening coverage to additional subjects, grade levels, and languages.
@@ 5 Future Work
As this work is exploratory, we do not anticipate any significant ethical risks associated with it. Moreover, one of our main goals in this research is to raise awareness of the education stakeholders as well as practitioners interested in the state of the art in AI in Education of the current capabilities of LLMs-as-tutors. Through transparent evaluation of such models, we aim to improve their interpretability, which we hope will help avoid potential future risks associated with a wider adoption of these models in education.

### v2

# Kseniia Petukhova and Ekaterina Kochmar

@@ Abstract
Numaan Naeemthanks: Equal contribution., Kaushal Kumar Maurya, Kseniia Petukhova and Ekaterina Kochmar Mohamed bin Zayed University of Artificial Intelligence, Abu Dhabi, UAE {Numaan.Naeem, kaushal.maurya, kseniia.petukhova, ekaterina.kochmar}@mbzuai.ac.ae
@@ 1 Introduction
The current version of our tool focuses on the pedagogical quality evaluation of tutor responses in the context of Student Mistake Remediation (SMR) Boaler (2013); Handa et al. (2023) in the mathematical domain, where tutors address errors or misconceptions that hinder students’ progress Wang et al. (2024a, b); Macina et al. (2023a). As the foundation for evaluation, we use the established taxonomy from Maurya et al. (2025), which is grounded in the learning sciences principles and which allows us to assess the quality of tutor responses along four key SMR dimensions Kochmar et al. (2025): (1) mistake identification, concerned with whether tutor’s response notifies the student of the committed mistake; (2) mistake location, focusing on whether the tutor clearly points to the erroneous part in the student’s solution; (3) providing guidance, evaluating the quality of the pedagogical guidance; and (4) actionability, assessing whether tutor’s response makes it clear what the student should do next. These dimensions are further outlined in Table 3, with an illustrative example provided in Figure 1.
The structure and implementation of the front- and backend of our tool are described in Section 3. Using MRBench dataset Maurya et al. (2025) and taking inspiration from the BEA 2025 shared task on AI tutor response evaluation Kochmar et al. (2025), we introduce a novel, efficient, and light-weight multi-task learning model that addresses the four dimensions of pedagogical quality evaluation ($\S$3.1). The outputs of the model, as well as those of an open-source (Prometeus Kim et al. (2024)) and a commercial (GPT-5) LLMs used as judges, are displayed using an interactive browser-based UI with helpful visualizations ($\S$3.3). The evaluation results ($\S$4) suggest that while our model achieves competitive results when evaluated against gold standard annotation, its outputs are also perceived by users to be at least as accurate as those of the commercial LLM-as-judge models, and the UI is considered informative and easy to use. We present and publicly release:
@@
  • An interactive UI available at https://demo-ai-tutor.vercel.app, which communicates the results and showcases the capabilities of AI tutors in an interpretable way, which we consider to be of interest to education stakeholders and the *ACL community at large. The demo tool can also be run locally with the user’s own data and models following the instructions provided in the GitHub repository.
@@ 2 Related Work
Over the years, the NLP community has seen significant advances in the development of publicly available toolkits for modeling and evaluation, including Hugging Face Transformers Wolf (2019), NLTK Bird (2006), and Scikit-learn Pedregosa et al. (2011). These toolkits have enhanced code reusability, enabling researchers to focus on developing more sophisticated models and metrics.
However, in the educational domain, especially in conversational dialogues, there remains a lack of robust tools to push research boundaries. Some progress has been made with frameworks like ConvoKit Chang et al. (2020), which facilitates the manipulation and analysis of general conversational data, and the social interactions embedded within. More recently, Edu-ConvoKit Wang and Demszky (2024) was developed for preprocessing, annotation, and analysis specifically for educational dialogues. While these toolkits contribute significantly to data handling and analysis, they fall short of addressing the evaluation of pedagogical quality in AI-driven educational systems. To the best of our knowledge, there is no toolkit that supports on-the-fly assessment of the pedagogical abilities of AI tutors. With AITutor-EvalKit, we aim to fill this critical gap by providing an open-source toolkit for the systematic evaluation of AI tutors. The toolkit integrates a popular taxonomy proposed by Maurya et al. (2025) into an automated evaluation framework. It is also designed to be extensible to additional evaluation aspects such as the ones proposed by Macina et al. (2025), who augment pedagogical assessment with measures of student expertise and understanding. By supporting and unifying such evaluation methods, AITutor-EvalKit aims to facilitate progress in this underexplored yet important research area.
Figure 2: AITutor-EvalKit pipeline: The backend module includes several model options to assess the pedagogical soundness of tutors’ responses, and the frontend presents evaluation outputs in an interactive user interface.
@@ 3 System and Demonstration Description
AITutor-EvalKit consists of two major modules: backend and frontend, with the pipeline illustrated in Figure 2. The backend module includes several models for tutor response evaluation, and frontend seamlessly integrates these evaluation outputs with an interactive, customizable, and flexible UI. Different audience groups can benefit from different functionalities of the toolkit: e.g., researchers and developers can use both modules, train their own automated models, use their own datasets, and launch the demo UI locally, while teachers, policymakers and other educational stakeholders can use the frontend module to understand the capabilities of LLMs-as-tutors and make decisions accordingly.
@@ 3.1 Backend Module: Evaluation Models
The backend module has two components: a specialized automated evaluation model and an evaluation pipeline for LLMs-as-judges. In this section, we provide details on the task, the training data (for the automated evaluation model), the test data, and the automated evaluation model itself, as well as the usage of an open-source LLM and a closed-source GPT model as judges.
3.2 Student Mistake Remediation Task
This task considers mathematical educational dialogues between a student and a tutor, where interactions are driven by student’s mistakes or confusions, and the AI tutor aims to remediate them through pedagogically appropriate responses. Formally, let the conversation history be $H=\{(T_{1},S_{1}),(T_{2},S_{2}),\dots,(T_{t},S_{t})\}$, where $T_{i}$ and $S_{i}$ denote the tutor’s and student’s $i$-th utterances, respectively. Let $S_{k}$, $k\in[1,\dots,t]$, denote the most recent student’s utterance containing a mistake or confusion; the tutor then produces $T_{t+1}$ to address it. The proposed toolkit evaluates the pedagogical quality of $T_{t+1}$ along eight dimensions defined by Maurya et al. (2025).
@@ 3.2.1 Dataset
As discussed in Section 1, we used MRBench dataset by Maurya et al. (2025), which has 491 dialogues (300 in the development set and 191 in the test set), each paired with seven LLM-generated tutor responses and one or two human tutor responses. Each response is annotated by human annotators for each of the four dimensions using the categories “Yes,” “To some extent,” and “No.” We also created a randomly selected demonstration set consisting of 10 dialogues, which is a subset of the test set. More details on the dimensions, annotation, and data statistics can be found in Appendix A.
3.2.2 Specialized Automated Model: LoMTL
This component implements a ternary classifier to evaluate tutor responses across the four considered dimensions. The models developed by participants in the BEA shared task provide a good starting point, as summarized by Kochmar et al. (2025). However, upon closer inspection, we found that several teams relied on closed-source models, fully fine-tuned and LoRA-tuned models, and some even used multiple models or ensembles for each dimension, which makes these approaches difficult to scale across all dimensions.
@@ 3.2.3 LLM-as-a-Judge Evaluation
This component provides functionality for evaluation using LLMs-as-judges. Following Maurya et al. (2025), we have selected Prometheus2 Kim et al. (2024) as the primary open-source LLM. However, the implementation of this module is flexible enough to support the family of Llama and other open-source causal LLMs for evaluation. Additionally, we have included an option to evaluate using the closed-source OpenAI GPT-5 model, although the implementation supports any model from the GPT family. Users must provide their OpenAI API key to run evaluations. We selected the above two models considering their open- and closed-source nature, as well as their human-like performance on public benchmarks Kim et al. (2024). However, the codebase is flexible enough to require only minimal adaptation to support other LLMs.
@@ 3.3 Frontend Module: Demo App
We present an interactive prototype built on top of MRBench, designed to evaluate the pedagogical abilities of AI tutors in educational dialogues. Users can explore the demo in two modes: (1) a Static Mode, where educators can access a deployed version containing 10 conversations from the MRBench test set evaluated by our fine-tuned model and share feedback on the helpfulness of tutor responses; and (2) a Customized Mode, where developers can run the interface locally to analyze their own datasets using either our LoMTL model or their own one, supported by the provided code (the exact step-by-step details are supplied in the official GitHub README file [^2] https://github.com/kaushal0494/AITutor-EvalKit/blob/main/README.md). The UI consists of three key modules – Automated Evaluation, LLM Evaluation, and Visualizer, that together enable exploration, analysis, and visualization of the key aspects of educational dialogues.
@@ 3.3.1 Automated Evaluation UI
The Automated Evaluation module presents results generated by our LoMTL model, assessing AI tutors across four pedagogical dimensions. As shown in Figure 6, users are first presented with an overview panel summarizing key statistics such as the number of topics, models, dimensions, and conversations. Users can evaluate a single tutor or compare two tutors side by side.
Single Tutor Evaluation: In this mode, users select a problem topic to view the complete student-tutor dialogue in the “Context” block (Figure 7). A drop-down menu allows selection of a tutor model, and the corresponding response is displayed in the “Tutor Response” block. Users can rate the usefulness of the response (Helpful, To Some Extent, or Not Helpful) and optionally view the ground truth solution for better context. Upon clicking “Get Auto-Evaluation Results,” the system generates performance evaluations across the chosen dimensions, with results downloadable in a PNG, JPG, or JSON format. The “Best Performance by Dimension” panel highlights the top tutor(s) for each dimension, helping users quickly identify their pedagogical strengths.
@@ 3.3.2 LLM Evaluation UI
The LLM Evaluation module enables advanced pedagogical assessment of AI tutor responses using LLMs as judges. It extends the functionality of the automated evaluation pipeline by leveraging LLMs to judge and compare tutor responses across four pedagogical dimensions. The UI supports three evaluation modes: single tutor evaluation, tutor model comparison, and LLM judge comparison. As shown in Figure 10, the overview panel summarizes available topics, conversations, tutor models, evaluation dimensions, and judge LLMs, providing a quick snapshot of the evaluation setup. Currently, two LLM judges are supported: GPT-5 and Prometheus-7B-v2.0 Kim et al. (2024).
@@ 3.3.3 Visualizer UI
The Visualizer module provides a high-level overview of the MRBench development set, using gold-standard annotations across four evaluation dimensions through intuitive visual analytics. Users are first presented with a “Dataset Overview” panel summarizing key statistics, including the number of conversations, tutor models, and evaluation dimensions. This module includes three main visualization panels: Tutor Performance Summary, Visualization Controls, and Dataset Visualization.
The Tutor Performance Summary panel presents average scores for each tutor model across all four dimensions, where “Yes,” “To some extent,” and “No” correspond to 1.0, 0.5, and 0.0, respectively (Figure 12). The Visualization Controls panel allows users to select specific tutors and dimensions to generate detailed visualizations (Figure 13). The Dataset Visualization panel then displays the results through spider and bar charts, where spider plots (the most informative visualizations from our perspective) summarize tutors’ strengths and weaknesses across dimensions (Figure 14), while bar charts show detailed score distributions for selected dimensions (Figure 15), along with averages for each response label.
@@ 4 Evaluation
To assess the toolkit and evaluation models, we first measure models’ performance using the metrics from the BEA 2025 shared task Kochmar et al. (2025) – accuracy and macro-F1, and then conduct a human evaluation study, in which participants assess both the prediction quality of the LoMTL evaluation model and the usability of our demo tool and UI.
@@ 4.1 Intrinsic Evaluation: Quantitative Analysis
For intrinsic evaluation, we compare our LoMTL model’s predictions on the test set from Kochmar et al. (2025) with the gold human annotations and also run GPT-5 and Prometheus2 on the same data. The average results across evaluation dimensions, presented in Table˜1, show that Prometheus2 performs substantially worse than both our model and GPT-5. Our model achieves the highest averaged accuracy and macro-F1 on the full test set, and it also performs competitively on the demonstration subset. While GPT-5 achieves a slightly higher averaged macro-F1 on the demonstration subset, the overall trend indicates that our model provides more reliable evaluations than Prometheus2 and performs on par with, or better than, GPT-5. A close inspection of the confusion matrix between human annotations and the LoMTL model shows strong overall agreement, with the majority of instances concentrated on the diagonal, particularly for clear Yes (3,106) and No (1,057) cases. However, the model exhibits systematic confusion on borderline instances and a mild tendency to over-predict Yes, especially when humans assign To some extent or No labels. Extended results and a related discussion, including average precision and recall scores, are provided in Table 7 in Appendix C.
@@
Prometheus2 0.47 0.41 0.41 0.34
GPT-5 0.66 0.58 0.66 0.59
Table 1: Accuracy and macro-F1 scores (averaged across Mistake Identification (MI), Mistake Location (ML), Providing Guidance (PG), and Actionability (AC)) for LoMTL, Prometheus2, and GPT-5 on the full test set from Kochmar et al. (2025) and on the demonstration set. Best results are shown in bold.
Performance across dimensions: The per-dimension results, presented in Table˜2, show that Prometheus2 performs poorly, while our model notably outperforms GPT-5 on the Mistake Identification and Actionability dimensions. However, it underperforms GPT-5 by 3 and 6 percentage points in terms of macro-F1 for Mistake Location and Providing Guidance, respectively. A similar trend is observed in the ten dialogues used for demonstration.
Model Full Test Set Demonstration Set
Accuracy Macro-F1 Accuracy Macro-F1
MI ML PG AC MI ML PG AC MI ML PG AC MI ML PG AC
@@ 4.2 Extrinsic Evaluation: Human Study
5 Conclusions and Future Work
This paper introduces the first open-access, open-source model for pedagogical quality evaluation of AI tutor responses, released under an MIT license. It is grounded on pedagogical principles and presents multiple evaluation choices including our proposed light-weight multi-tasking LoMTL model. The toolkit is highly customizable, allowing researchers and practitioners to extend it to diverse educational contexts and dialogues across domains, and it can be easily set up and run locally. In addition, we provide an interactive web-based UI that offers interpretable evaluations of the pedagogical capabilities of state-of-the-art LLMs acting as AI tutors for education stakeholders, policymakers, and non-technical audience.
Future work will extend the toolkit by (1) integrating new user dialogues and LLMs in the frontend module to generate real-time responses and evaluations; (2) expanding the evaluation dimensions; (3) enabling new data upload option in UI; and (4) broadening coverage to additional subjects, grade levels, and languages.
@@ 5 Conclusions and Future Work
As this work is exploratory, we do not anticipate any significant ethical risks associated with it. Moreover, one of our main goals in this research is to raise awareness of the education stakeholders as well as practitioners interested in the state of the art in AI in Education and the current capabilities of LLMs-as-tutors. Through transparent evaluation of such models, we aim to improve their interpretability, which we hope will help avoid potential future risks associated with a wider adoption of these models in education.
This work uses the MRBench dataset Maurya et al. (2025), which integrates the MathDial Macina et al. (2023a) and Bridge Wang et al. (2024a) datasets. In these datasets, the identities of the tutors are not revealed, while the student profiles are either synthetically created or anonymized. As a result, we do not anticipate any direct ethical risks associated with the datasets used.